# Estimating Correctness Without Oracles in LLM-Based Code Generation

### v1

@@ Appendix A Proof of Pointwise Incoherence Inequality
Appendix B Estimation of Incoherence and Error
B.1 Monte Carlo Estimation of Pointwise Error
@@
Output: ¯$\bar{\operatorname{\mathcal{E}}}_{\mathsf{Gen}}(d)$ estimate such that |¯$|\bar{\operatorname{\mathcal{E}}}_{\mathsf{Gen}}(d)-\operatorname{\mathcal{E}}_{\mathsf{Gen}}(d)|\leq\epsilon$ with probability $\geq 1-\delta$
@@
5:     $e_{i}\leftarrow\mathbb{I}_{\llbracket\pi\rrbracket(x)\neq f^{\ast}_{d}(x)}$
@@
7:  return ¯$\bar{\operatorname{\mathcal{E}}}_{\mathsf{Gen}}(d)=\frac{1}{N}\sum_{i=1}^{N}e_{i}$
@@
(|¯$\mathbb{P}(|\bar{\operatorname{\mathcal{E}}}_{\mathsf{Gen}}(d)-\operatorname{\mathcal{E}}_{\mathsf{Gen}}(d)|\geq\epsilon)\leq 2\exp(-2N\epsilon^{2}).$
@@
B.2 Monte Carlo Estimation of Pointwise Incoherence
@@
Output: ¯$\bar{\operatorname{\mathcal{I}}}_{\mathsf{Gen}}(d)$ estimate such that |¯$|\bar{\operatorname{\mathcal{I}}}_{\mathsf{Gen}}(d)-\operatorname{\mathcal{I}}_{\mathsf{Gen}}(d)|\leq\epsilon$ with probability $\geq 1-\delta$
@@
5:     $d_{i}\leftarrow\mathbb{I}_{\llbracket\pi_{1}\rrbracket(x)\neq\llbracket\pi_{2}\rrbracket(x)}$
@@
7:  return ¯$\bar{\operatorname{\mathcal{I}}}_{\mathsf{Gen}}(d)=\frac{1}{N}\sum_{i=1}^{N}d_{i}$
@@
(|¯$\mathbb{P}(|\bar{\operatorname{\mathcal{I}}}_{\mathsf{Gen}}(d)-\operatorname{\mathcal{I}}_{\mathsf{Gen}}(d)|\geq\epsilon)\leq 2\exp(-2N\epsilon^{2}).$
@@
B.3 PAC Detection of Nonzero Incoherence
@@
Appendix C Functional Incoherence
@@
C.1 Functional Incoherence as a Lower Bound on Error
Theorem C.1 (Functional Incoherence Inequality).
@@ Discussion.
Table 5: Interpretation of Spearman’s $\rho$, based on thresholds from Schober, Boer, and Schwarte (2018).

### v2

@@ Appendix B Proof of Pointwise Incoherence Inequality
Appendix C Estimation of Incoherence and Error
C.1 Monte Carlo Estimation of Pointwise Error
@@
Output: $\bar{\operatorname{\mathcal{E}}}_{\mathsf{Gen}}(d)$ estimate such that $|\bar{\operatorname{\mathcal{E}}}_{\mathsf{Gen}}(d)-\operatorname{\mathcal{E}}_{\mathsf{Gen}}(d)|\leq\epsilon$ with probability $\geq 1-\delta$
@@
5:  $e_{i}\leftarrow\mathbb{I}\bigl(\llbracket\pi\rrbracket(x)\neq f^{\ast}_{d}(x)\bigr)$
@@
7:return $\bar{\operatorname{\mathcal{E}}}_{\mathsf{Gen}}(d)=\frac{1}{N}\sum_{i=1}^{N}e_{i}$
@@
$\mathbb{P}(|\bar{\operatorname{\mathcal{E}}}_{\mathsf{Gen}}(d)-\operatorname{\mathcal{E}}_{\mathsf{Gen}}(d)|\geq\epsilon)\leq 2\exp(-2N\epsilon^{2}).$
@@
C.2 Monte Carlo Estimation of Pointwise Incoherence
@@
Output: $\bar{\operatorname{\mathcal{I}}}_{\mathsf{Gen}}(d)$ estimate such that $|\bar{\operatorname{\mathcal{I}}}_{\mathsf{Gen}}(d)-\operatorname{\mathcal{I}}_{\mathsf{Gen}}(d)|\leq\epsilon$ with probability $\geq 1-\delta$
@@
5:  $d_{i}\leftarrow\mathbb{I}\bigl(\llbracket\pi_{1}\rrbracket(x)\neq\llbracket\pi_{2}\rrbracket(x)\bigr)$
@@
7:return $\bar{\operatorname{\mathcal{I}}}_{\mathsf{Gen}}(d)=\frac{1}{N}\sum_{i=1}^{N}d_{i}$
@@
$\mathbb{P}(|\bar{\operatorname{\mathcal{I}}}_{\mathsf{Gen}}(d)-\operatorname{\mathcal{I}}_{\mathsf{Gen}}(d)|\geq\epsilon)\leq 2\exp(-2N\epsilon^{2}).$
@@
C.3 PAC Detection of Nonzero Incoherence
@@
Appendix D Functional Incoherence
@@
D.1 Functional Incoherence as a Lower Bound on Error
Theorem D.1 (Functional Incoherence Inequality).
@@ Discussion.
Table 5: Interpretation of Spearman’s $\rho$, based on thresholds from schober2018correlation.